find mirr
Homework Due 10/29 I= 10% ? ? - find me -1000 you -200 2000 100.
Answer #1

MIRR = ( FVc / PVfc )1/n -1

where, FVc is the  future value of positive cash flows

and PVfc is the present value of negative cash flows

CF0 = -1000
CF1 = 800
CF2 = -200
CF3 = 2000
CF4 = 100

Interest Rate = i = 10%

PVfc = 1000 + 200/(1+0.10)2 = $1165.29

FVc = 800*(1+0.10)3 + 2000*(1+0.10) + 100 = $3364.8

Hence, MIRR = (3364.8/1165.29)1/4 - 1 = 0.3036 = 30.36%